y varies directly as x and inversely as the square of z. y equals 2=20 when x equals =18 and z equals =3. Find y when x equals =2 and z equals =8.
Answer by Boreal(15235)

y=kx/z^2
substitute and assume y=20, not 2
20=k*18/9=2k
k=10
now put that into original equation
y=10*2/64=20/64, or 5/16